In a spherical capacitor with plate radii a and b, the PD between them is V. The electric field between the platesnat a radial distance r from the centre is?

Grade:12

4 Answers

Arun Kumar IIT Delhi
9 years ago
Hi Mostafijur,

Let the inner sphere has a charge +Q
outer sphere -Q
then
between the sphere we have

Thanks & Regards

Arun Kumar

IIT Delhi

Askiitians Faculty

Yash Chourasiya
3 years ago
Hello Student

Let potential of the plate of radius a
Q = CV
V = Q/C
E = KQ/r2
V = KQ(b−a)/ab
Q = Vab/K(b−a)
E = (K/r2)(Vab/K(b−a)) = abV/r2(b−a)

I hope this answer will help you.
Mansi
one year ago
As v= kq[1/a-1/b]kq=abv/b-a
Mansi
one year ago
As v= kq[1/a-1/b]kq=abv/b-aE=kq/r^2=abv/r^2(b-a)Hope this answer helps you a lot !Thanks and RegardsMansi